Nico is saving money for his college education. He invests some money at 9​%, and ​$1400 less than that amount at 6%. The investments produced a total of ​$201 interest in 1 yr. How much did he invest at each​ rate?
Answer by ankor@dixie-net.com(22740)

Nico is saving money for his college education.
He invests some money at 9​%, and ​$1400 less than that amount at 6%.
The investments produced a total of ​$201 interest in 1 yr.
How much did he invest at each​ rate?
:
Let x = amt invested at 9$
then
(x-1400) = amt invested at 6%
:
Total interest equation
.09x + .06(x-1400) = 201
.09x + .06x - 84 = 201
.15x = 201 + 84
.15x = 285
x = 285/.15
x = $1900 invested at 9%
then
1900 - 1400 = $500 invested at 6%
:
:
Check this find the actual interest at each
.09(1900) = 171
.06(500) = 30
----------------
total int: $201